OpenStudy (anonymous):

point (-2,4) and (-1,1) lie on the same line. which of the following equations represent the line and why? A.y-1=-1/3(x+1) B. y-1=-3(x+1) C.y-1=-1/3(x-1) D.y-1=-3(x-1)

OpenStudy (anonymous):

this arent test nor homework its just old study guides i got from a random website to help me study for my EOCT exams coming up in a week or so

OpenStudy (anonymous):

Simply this and you should get your answer: \[\frac{ y-4 }{ x - (-2) } = \frac{ 1 - 4 }{ -1 - (-2) }\]

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

In hindsight: \[{ y - 1 } = \frac{ 4 - 1 }{ -2 - (-1) } \times (x - (-1))\] Would be faster

OpenStudy (anonymous):

kk hold on

OpenStudy (anonymous):

is it A

OpenStudy (anonymous):

@MonaCraft